QuinnLoves:

Find the GCF of 56 and 96. Show your work.

QuinnLoves:

I know the answer is 8 but, I don't know how to show my work.

QuinnLoves:

Oh no youve done enoguh lol

QuinnLoves:

Your too helpfull XD

QuestionCoveBot:

Alright. Just like the last question. 56 is not a prime number. \[2 * 2 * 2 * 7 => 2^3 * 7^1\]96 is also not a prime number. \[2 * 2 * 2 * 2 * 2 * 3 => 2^5 * 3^1\]New list is... 2,2,2,2,2,3,7 Multiply each one... \[LCM = 2 * 2 * 2 * 2 * 2 * 3 * 7 = 672\]Exponential form... \[LCM = 2^5 * 3^1 * 7^1 = 672\]LCM = 672 So therefore, LCM(56, 96) = 672, not 8.

QuinnLoves:

WHYYYYY, Your to helpful

QuestionCoveBot:

Just trying to help people. :) Way I am. Your welcome.

QuinnLoves:

:)

Babbbygirlnini:

Factors of 56: 1, 2, 4, 7, 8, 14, 28, 56 Factors of 96: 1, 2, 3, 4, 6, 8, 12, 16, 24, 32, 48, 96 GCF(56, 96) = 8

QuestionCoveBot:

I finally noticed it was GCF...

QuestionCoveBot:

Not LCM... Sorry. e.e

QuestionCoveBot:

Let me take 5 minutes to redo this. :/

Florisalreadytaken:

redo what? that other person did it already haha!

QuestionCoveBot:

Yeah, right. ._.

QuestionCoveBot:

Sorry @QuinnLoves

QuinnLoves:

Yea, but different numbers

QuinnLoves:

NP AT ALL

QuinnLoves:

If anything, I should be saying Sorry To YOU. I made ou do TWO of these XD

QuinnLoves:

I promise I wont ask more lol

QuestionCoveBot:

It is fine.

QuinnLoves:

B)